Based on the inverse square law, at distance d = 5 m: (a) what is the total area over which energy is dispersed? (b) what is the fraction of the total area represented by the total dispersal area compared to d = 1? 


1
Expert's answer
2021-10-12T12:33:17-0400

a)

sphere surface area:

"A=4\pi d^2=4\pi\cdot 5^2=100\pi=314" m2


b)

the fraction:


"\frac{4\pi\cdot 1^2}{4\pi\cdot 5^2}=\frac{1}{25}"